### v2

# Evaluating NeRFs for 3D Plant Geometry Reconstruction in Field Conditions

Abstract.
We evaluate different Neural Radiance Fields (NeRFs) techniques for reconstructing (3D) plants in varied environments, from indoor settings to outdoor fields. Traditional techniques often struggle to capture the complex details of plants, which is crucial for botanical and agricultural understanding. We evaluate three scenarios with increasing complexity and compare the results with the point cloud obtained using LiDAR as ground truth data. In the most realistic field scenario, the NeRF models achieve a 74.65% F1 score with 30 minutes of training on the GPU, highlighting the efficiency and accuracy of NeRFs in challenging environments. These findings not only demonstrate the potential of NeRF in detailed and realistic 3D plant modeling but also suggest practical approaches for enhancing the speed and efficiency of the 3D reconstruction process.
Neural Radiance Fields (NeRFs); 3D Geometry Reconstruction; Field Conditions
Figure 1. Using NeRFs for 3D reconstruction of plants in field conditions.
1. Introduction
In recent years, the importance of reconstructing 3D geometry has emerged as a critical area within plant sciences. As global challenges in agriculture and botany become increasingly complex [Pereira, 2017], gaining a detailed understanding of plant structures has become essential. This goes beyond mere visual representation; capturing the intricate details of plant morphology provides valuable insights into their growth, responses to environmental stressors, and physiological processes [Kumar et al., 2012, Paturkar et al., 2021a]. NeRF emerges as a pivotal innovation in the domain of 3D geometry reconstruction.
At its core, NeRF (Neural Radiance Fields) represents a groundbreaking technique that utilizes the capabilities of deep learning to synthesize continuous 3D scenes by modeling the volumetric scene function [Mildenhall et al., 2021]. This approach significantly advances beyond traditional methods such as structure from motion (SfM) [Eltner and Sofia, 2020] and Multi View Stereo (MVS) [Chen et al., 2019], which rely on the analysis of discrete 2D pixels, thus marking a new era in the field of 3D model reconstruction.Unlike its predecessors, NeRF enables the rendering of photorealistic scenes from any viewpoint using a set of 2D images, without necessitating explicit 3D geometry or depth maps. This is primarily due to NeRF’s use of implicit representation models, in contrast to the explicit representations like point clouds in Structure from Motion (SfM) and voxel grids in Multi View Stereo (MVS), which inherently limit the quality of the reconstruction. The implicit representation utilized by NeRF is resolution invariant, allowing for significantly more detailed and granular modeling without the constraints of resolution-dependent methods. This innovation not only enhances the reconstruction of complex scenes but also holds the promise of revolutionizing the 3D modeling of plant structures by providing an unprecedented level of detail and granularity. The versatility and rapid adoption of NeRF as a state-of-the-art technique underscore its significance, with applications ranging from virtual reality [Deng et al., 2022] to architectural reconstructions [Tancik et al., 2022]. Particularly in botany and plant research, NeRF’s ability to capture the finest details offers potential for groundbreaking insights into plant structures, establishing it as a key tool in the advancement of modern botanical studies. However, translating this potential into practical applications presents its own set of obstacles.
1.1. Challenges in 3D Plant Modeling
Contemporary 3D modeling techniques for plant structures face significant challenges when attempting to capture the minute details inherent in plants [Kumar et al., 2012]. The complexity of plants, from their delicate leaf venation [Lu et al., 2009] to intricate branching patterns [Evers, 2011], necessitates models that encompass these specific details. A summary of these is given in Table 1
1.1.1. Technical Challenges
Traditional methodologies like photogrammetry are adept at large-scale reconstructions, but often overlook these subtle intricacies [Paturkar et al., 2019, Wu et al., 2020, Wang et al., 2022]. Utilizing tools such as FARO 3D LiDAR scanners have been adopted for 3D reconstructions of plants like maize [Wu et al., 2020] and tomatoes [Wang et al., 2022]. However, to capture every detail, data from multiple angles is essential. This poses a challenge since scanning from each side using devices like FARO 3D LiDAR scanners is time-consuming, limiting the number of angles that can be efficiently captured. Due to the limited poses, this approach does not scale well to capture minute details in large scenes, and consequently, some desired details might be missed in the final model. Andújar et al. have emphasized that, even with advanced sensors, there are gaps in detailed reconstruction and the preservation of essential information for intricate applications [Andújar et al., 2018]. They also point out that while devices such as the MultiSense S7 from Carnegie Robotics combine lasers, depth cameras, and stereo vision to offer commendable results, the high acquisition costs can be prohibitive. The complexity extends beyond static to dynamic realms.
In addition to the aforementioned challenges, the dynamic nature of flexible objects like plants and environmental dynamism introduce an added dimension of complexity. Plants, unlike static entities, undergo growth, exhibit movement in reaction to environmental stimuli, and demonstrate both diurnal and seasonal variations. The environmental dynamism, coupled with plant behavior, further complicates modeling efforts. Paturkar et al.’s comprehensive investigation underscores that this dynamism inherently complicates the attainment of precise 3D models. Factors such as persistent growth, environmental dynamism, and external perturbations, notably in windy scenarios, jeopardize the consistency of data acquisition during imaging processes [Paturkar et al., 2019, Lu, 2023, Paturkar et al., 2021b]. This complexity necessitates innovative solutions in 3D modeling and data processing.
Furthermore, an array of 3D modeling approaches mandate rigorous [Setkowicz, 2014], an exercise that necessitates vast computational capacities and specialized knowledge. This post-processing phase, being resource and time-intensive, often becomes a constraining factor in projects, particularly when modifications or refinements are imperative based on novel data or feedback. As emphasized in Liénard et al. [2016], discrepancies during the post-processing phase in UAV imagery-centric 3D scene reconstruction might culminate in costly and irreversible repercussions. This necessitates a reevaluation of the cost-benefit ratio inherent in these technologies.
1.1.2. Accessibility Challenges
These factors indicate that the challenges in capturing every facet of plant structures remain, even when employing sophisticated sensors. Financial implications further exacerbate these challenges. Although techniques like terrasential LiDAR scanners provide superior accuracy, their exorbitant costs often render them inaccessible to a significant portion of researchers. Tang et al. delineate that the substantial financial commitment associated with such advanced equipment, combined with the specialized expertise requisite for its operation, significantly limits their adoption within both academic and enthusiast domains [Tang et al., 2022]. Recognizing these barriers, our study proposes an innovative solution.
Table 1. Challenges in 3D Modeling Techniques for Plant Structures
Technical Challenges Accessibility Challenges
Detail limitations in Photogrammetry High acquisition costs
Data acquisition from multiple angles Limited equipment accessibility
Dynamic nature of plants Expertise for operation
Environmental dynamism
Rigorous post-processing
1.2. Objectives and Significance of the Study
This study focuses on a detailed evaluation of advanced Neural Radiance Fields (NeRF) methodologies to assess their applicability and effectiveness in high-resolution 3D modeling of plant structures. Traditional 3D modeling techniques often fall short when it comes to accurately capturing the complex and dynamic morphology of plants [Nguyen et al., 2015]. Our research aims to explore how NeRF can address these limitations, offering a more nuanced and detailed approach to modeling plant structures. Bridging this gap, our study integrates cutting-edge NeRF techniques with mobile technology advancements.
A pivotal aspect of our methodology is the use of mobile technology for data acquisition. By utilizing the widespread availability and sophisticated capabilities of modern smartphones, we aim to make high-quality 3D data collection more accessible and cost-effective. This approach, combined with the NeRF framework’s ability to process a wide variety of image datasets, is expected to revolutionize 3D plant modeling practices, enhancing both inclusivity and efficiency. Next, we delve into the specifics of our research.
An essential part of our study involves a comparative analysis of different NeRF implementations to determine the most effective configurations for specific modeling needs. This includes assessing detail fidelity, computational efficiency, and the ability to adapt to changes in environmental conditions. Such comparative analysis is crucial for establishing benchmarks for NeRF’s current capabilities and for identifying opportunities for future technological improvements. We summarize our contributions as follows:
  1. (1)
    A dataset collection encompassing a wide range of plant scenarios for reconstruction purposes consisting of images, camera poses, and ground truth TLS scans.
  2. (2)
    An evaluation of state-of-the-art NeRF techniques across seven different metrics, offering insights for further research.
  3. (3)
    The development of an end-to-end evaluation framework designed to reassess future reconstruction technologies, specifically for 3D modeling in plant sciences.
Furthermore, a comprehensive review of related works and NeRF techniques mentioned in the literature is provided in Table 2, setting the stage for our investigation.
Table 2. Techniques in Literature.
Used original paper implementation.
Used implementation from Framework (NeRFStudio/SDFStudio).
Instant-NGP NeRFacto Mip-NeRF NeRFs Additional Techniques
Towards a Robust Framework for NeRF Evaluation[Azzarelli et al., 2023]
$\times$ $\times$
Analyzing the Internals of Neural Radiance Fields[Radl et al., 2023]
$\times$ $\times$
SteerNeRF: Accelerating NeRF Rendering via Smooth Viewpoint Trajectory[Li et al., 2023]
$\times$ $\times$ NSVF, KiloNeRF, PlenOctree, DIVeR
A Critical Analysis of NeRF-Based 3D Reconstruction[Remondino et al., 2023]
$\times$ $\times$ Tensorf, Mono-Neus, Neus-Facto, MonoSDF, VolSDF, NeuS, UniSurf
Few-Shot Photogrametry: NeRF vs. MVS-SfM in Cultural Heritage Documentation[Balloni et al., 2023]
$\times$ $\times$ $\times$ $\times$
2. Materials and Methods
2.1. Field scenarios
In order to evaluate the efficacy of NeRF techniques across varying complexities and conditions, three scenarios are considered:
  1. (1)
    Single Corn Plant Indoor: This serves as the simplest test case. A solitary corn plant is placed in a controlled indoor environment. The lighting, background, and other environmental factors are kept constant. The objective is to assess the basic capabilities of NeRF in reconstructing an individual plant structure [Jignasu et al., 2023].
  2. (2)
    Multiple Corn Plants Indoor: In this case, more than one corn plant is situated in an indoor setting. The increased complexity due to multiple plants poses a greater challenge for the 3D reconstruction. Inter-plant occlusions and varying plant orientations add an additional layer of complexity.
  3. (3)
    Multiple Corn Plants in a Field with Other Plants: This scenario represents a real-world agricultural field, where corn plants are interspersed with other types of plants. The added complexity due to variable lighting, wind, and other dynamic environmental conditions tests the robustness of the NeRF technology.
2.2. Data Collection
In this experiment, our focus is exclusively on corn plants. We examine three distinct scenarios, elaborated subsequently, spanning from controlled indoor settings to dynamic outdoor ones. In the first scenario, we place a single corn plant in an indoor setting. In the second scenario, we place three corn plants side by side in indoor settings. In the third scenario, in the dynamic outdoor conditions of an experimental field at Iowa State University, we selected a row plot of corn plants planted at approximately 0.2 m distance, approximately at the V12 stage. The leaves between two neighboring plants are overlapping. Transitioning from the specific experimental setups, we now detail the data collection methods utilized in these scenarios.
Our training dataset for NeRF is sourced from RGB images and LiDAR data captured using a mobile phone, with the RGB images aiding in the 3D reconstruction of the plants and the LiDAR exclusively for pose capture. For all three scenarios, data is captured using an iPhone 13 Pro featuring 4K resolution. The device is held at a constant height while circling the plant to ensure consistent capture angles. The data collection process utilizes the Polycam app, with a time of approximately 2.5 minutes for scenario 3 (multiple plants in the outdoor setting) and around 1 minute for scenario 1 (single plant in the indoor setting). To establish accurate ground truth, we utilized high-definition terrestrial LiDAR scans using the Faro® Focus S350 Scanner. The scanner boasts an angular resolution of 0.011 degrees, equating to a 1.5 mm point spacing over a 10 m scanning range. With the capacity to acquire point clouds of up to 700 million points (MP) at 1 million points per second. Additionally, the scanner includes a built-in RGB camera that captures 360-degree images once the scanning process is complete.
Both in indoor and outdoor settings, we scan the plants from four (for the single plant) to six (for multiple plants) locations around the plant(s), at a height of 1.5 m and a distance of 1.5 m from the plant(s). To reduce the movement of the leaves during scanning, in indoor settings, we ensure that there is no airflow around the plants, and in outdoor settings, we waited for a suitable time when there was negligible wind flow (August 31, 2023, at 8:30 a.m.). Each scan required approximately 2 minutes and 24 seconds, totaling a capture time of around 18 minutes in outdoor settings, including manually moving the scanner around the plot. The six scans were processed in SCENE® software to add RGB color data to the point clouds, followed by the registration of the clouds by minimizing cloud-to-cloud distance and top view distance. Afterward, we cropped out the area of interest from the registered point cloud, removed duplicate points, and reduced noise using statistical outlier removal based on global and local point-to-point distance distributions. This process resulted in the point cloud having an average resolution of about 7 mm.
This experimental setup enables the NeRF algorithm to work on a range of complexities, from controlled environments to dynamic, real-world conditions.
(a) Frame 1
(b) Frame 2
(c) Frame 3
(d) Frame 4
Figure 2. Raw Images Provided to NeRFs - Scenario I
(a) Frame 1
(b) Frame 2
(c) Frame 3
(d) Frame 4
Figure 3. Raw Images Provided to NeRFs - Scenario II
(a) Frame 1
(b) Frame 2
(c) Frame 3
(d) Frame 4
Figure 4. Raw Images Provided to NeRFs - Scenario III
2.3. Camera Pose, NeRF Training, and Pointcloud Exports
Camera pose estimation is a crucial second step, typically achieved through a Structure from Motion (SfM) pipeline such as colmap. This process is essential for obtaining accurate 3D structures from sequences of images by determining correspondences between feature points and by using sequential matching, especially effective since our dataset comprises video frames.
In our workflow, we incorporate cutting-edge NeRF models for their rapid prototyping capabilities, which are critical in plant sciences to enable large scale reconstructions. Specifically, we employ Instant-NGP [Müller et al., 2022], Nerfacto [Tancik et al., 2023], and TensoRF [Chen et al., 2022], trained on an Nvidia A100 GPU with 80GB of memory. These models are at the forefront of efficiency, with vanilla NeRF requiring approximately 50 times longer to achieve comparable results. Notably, Instant-NGP introduces a small neural network complemented by a multiresolution hash table, optimizing the number of operations required for training and rendering [Müller et al., 2022]. TensoRF, on the other hand, conceptualizes the radiance field as a 4D tensor and applies tensor decomposition to achieve better rendering quality and faster reconstruction times compared to the traditional NeRF approach [Chen et al., 2022].
Nerfacto synthesizes various techniques from multiple contributions in the field, such as the Multilayer Perceptron (MLP) adapted from Instant-NGP, and the Proposal Network Sampler from MipNeRF-360 [Barron et al., 2022], as depicted in the attached image. Post-training, the models are converted into pointclouds with a million points each. The effectiveness of these reconstructions is then rigorously assessed as outlined in the next section.
2.4. Evaluation
2.4.1. Pipline
We reconstruct the scene and capture point clouds using a FARO scan for ground truth. To perform a one-to-one comparison between the NeRF-based reconstruction and ground truth, alignment is crucial. Our alignment and evaluation methodology is adapted from Knapitsch et al. [Knapitsch et al., 2017]. In their work, they evaluate different pipelines and use colmmap as a ’arbitrary reference’ frame. However, in our case, all the NeRFs use colmap in their pipeline so the reference frame and reconstruction frame becomes the same. This simplifies the steps and the updated steps, used in our paper are given below:
  1. (1)
    Preliminary Camera Trajectory Alignment: The NeRF-reconstructed point cloud is manually aligned with the ground truth using point-based alignment. Four corresponding points are selected in both point clouds to compute an initial transformation matrix. This matrix aligns the camera poses, providing initial scale and orientation estimates. This initial Coarse-grained alignment step paves the way for more detailed alignment procedures.
  2. (2)
    Cropping: Each ground-truth model has a manually-defined bounding volume, outlining the evaluation region for reconstructions.
  3. (3)
    ICP Registration: Drawing inspiration from the iterative refinement process detailed by Besl et al. [Besl and McKay, 1992] and further refined by Zhang et al. [Zhang, 1994], we adopt a three-stage approach as introduced by Knapitsch et al. [Knapitsch et al., 2017] for our initial registration framework. The process begins with a specified voxel size and an associated threshold for the initial registration. In the next iteration, the transformation result from the previous step is used as a starting point, with the voxel size reduced by half to achieve finer detail in the registration. The third stage aims to refine the alignment further by returning to the original voxel size and adjusting the threshold to facilitate convergence at each stage. This multi-scale strategy is designed to capture both coarse and fine details, thereby improving the accuracy and precision of the model alignment. However, in our adaptation for plant structure reconstruction, we diverged from Knapitsch et al. by maintaining the iterative process within a single stage rather than expanding across multiple stages. We found that increasing the iteration count tenfold, rather than the number of stages, prevented the registration process from collapsing [Billings et al., 2015].
2.4.2. Metrics
To assess the similarity between the ground truth (obtained from TLS) and the reconstructed 3D pointcloud, the following metrics are employed:
  1. (1)
    Precision/Accuracy. Given a reconstructed point set $\mathcal{R}$ and a ground truth set $\mathcal{G}$, the precision metric $P(d)$ assesses the proximity of points in $\mathcal{R}$ to $\mathcal{G}$ within a distance threshold $d$. Mathematically, it is formulated as:
    $P(d)=\frac{100}{|\mathcal{R}|}\sum_{\mathbf{r}\in\mathcal{R}}\mathbb{I}\left(\min_{\mathbf{g}\in\mathcal{G}}\|\mathbf{r}-\mathbf{g}\|<d\right),$
    where $\mathbb{I}(\cdot)$ is an indicator function.
  2. (2)
    Recall/Completeness. Conversely, the recall metric $R(d)$ quantifies how well the reconstruction $\mathcal{R}$ encompasses the points in the ground truth $\mathcal{G}$ for a given distance threshold $d$. It is defined as:
    $R(d)=\frac{100}{|\mathcal{G}|}\sum_{\mathbf{g}\in\mathcal{G}}\mathbb{I}\left(\min_{\mathbf{r}\in\mathcal{R}}\|\mathbf{g}-\mathbf{r}\|<d\right).$
    Both the above two metrics are extensively utilized in recent studies. [Balloni et al., 2023, Mazzacca et al., 2023].
  3. (3)
    F-score. The F-score, denoted as $F(d)$, serves as a harmonic summary measure that encapsulates both the precision $P(d)$ and recall $R(d)$ for a given distance threshold $d$. It is specifically designed to penalize extreme imbalances between $P(d)$ and $R(d)$. Mathematically, it can be expressed as:
    $F(d)=\frac{2\times P(d)\times R(d)}{P(d)+R(d)}.$
    The harmonic nature of the F-score ensures that if either $P(d)$ or $R(d)$ approaches zero, the F-score will also tend towards zero, providing a more robust summary statistic than the arithmetic mean. The value of $d$ was considered 0.005 for first two scenarios and 0.01 for the last scenario.
  4. (4)
    Traditional metrics such as, SSIM [Wang et al., 2004], PSNR [Hore and Ziou, 2010], LPIPS [Zhang et al., 2018] are used to evaluate the quality of the rendered images. These metrics do not need the 3D ground truth and are widely used in literature [Xu et al., 2022, Zhang et al., 2021] for evaluation.
(a) Ground Truth (LiDAR)
(b) Reconstruction (NeRF)
(c) Precision
(d) Recall
Figure 5. Visualization for Metrics
2.4.3. Visualization Color Interpretation
The color-coded visualizations employed in our research provide an intuitive understanding of spatial relationships within the 3D reconstructed plant structures. The color gradients signify varying distances from a specific reference point, enabling a comprehensive assessment of depth and intricacy. Specifically:
  1. (1)
    Grey: Represents regions of the plant structure that are closer or have the least distance to a particular reference.
  2. (2)
    Red: Depicts intermediate distances of points from the reference pointcloud within the plant structure.
  3. (3)
    Black: Highlights the furthest points or regions within the reconstructed structure.
3. Results
The results from all the scenarios are provided below:
3.1. Scenario I - Single Plants Indoors
Figure 6. Camera Pose Estimations for Scenario-I
The estimated camera poses from colmap are visualized in Fig. 6. The results from our evaluation of three state-of-the-art NeRF techniques for 3D plant structure reconstruction are illustrated in the Fig. 7.
(a) PSNR
(b) SSIM
(c) LPIPS
(d) Precision
(e) Recall
(f) F1 Score
Number of Iterations
  Instant-NGP    NeRFacto    TensoRF
Figure 7. Combined Comparison of 2D Image Quality (top) and 3D Geometry (bottom) Metrics for Scenario-I
Precision:
Across all models, precision generally increases with the number of iterations. Instant-NGP shows a significant leap in precision from 100 to 5000 iterations (0.29 to 21.93), indicating a drastic improvement in the accuracy of reconstructed points relative to the ground truth. NeRFacto demonstrates a more consistent and steep rise in precision, reaching a peak of 73.57 at 30000 iterations, which surpasses Instant-NGP’s best precision. TensoRF, however, shows a relatively modest increase in precision, suggesting its limited capability in accurately capturing fine details compared to the other two models. It is further illustrated in Fig. 8.
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Figure 8. Precision of 3D reconstruction of NeRF techniques
Recall:
The recall metric follows a similar trend, with Instant-NGP and NeRFacto showing substantial increases with more iterations, indicating an enhanced ability to encompass points from the ground truth. Notably, NeRFacto achieves remarkably high recall values (over 90) at higher iterations (10000 and above), suggesting its superiority in completeness of reconstruction. TensoRF’s recall values are significantly lower, suggesting that it may miss more details from the ground truth compared to the other models. It is further illustrated in Fig. 9.
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Figure 9. Recall of 3D reconstruction of NeRF techniques
F1 Score:
The F1 score, balancing precision and recall, highlights NeRFacto as the most balanced model, especially at higher iterations, with scores above 80. Instant-NGP shows a significant improvement in F1 scores as iterations increase, but it doesn’t reach the heights of NeRFacto. TensoRF lags behind in this metric, indicating a less balanced performance between precision and recall.
PSNR:
The Peak Signal-to-Noise Ratio (PSNR) reflects the quality of rendered images. In this metric, Instant-NGP and NeRFacto show a gradual increase in PSNR with more iterations, suggesting improved image quality. TensoRF’s PSNR values are lower, indicating potentially lower image quality throughout its iterations.
SSIM:
The Structural Similarity Index (SSIM) is another measure of image quality, assessing the perceived change in structural information. Here, NeRFacto and Instant-NGP both show a steady increase in SSIM with more iterations, with NeRFacto achieving slightly higher scores, suggesting better preservation of structural information in its renderings. TensoRF, again, shows relatively lower SSIM scores.
LPIPS:
The Lower Perceptual Image Patch Similarity (LPIPS) metric indicates perceived image similarity, with lower values being better. NeRFacto and Instant-NGP both show a significant decrease in LPIPS with more iterations, indicating improved perceptual similarity to the ground truth. TensoRF’s LPIPS values are consistently higher, suggesting lower perceptual similarity.
Computation Time:
Time efficiency is a crucial factor, especially for practical applications. Instant-NGP demonstrates a relatively balanced approach between efficiency and performance, with time increments correlating reasonably with the increase in iterations. However, it becomes significantly time-consuming at very high iterations (20000 and 30000). NeRFacto, while showing superior performance in many metrics, demands considerably more time, especially at higher iterations, which could be a limiting factor in time-sensitive scenarios. TensoRF, despite its lower performance in other metrics, maintains a more consistent time efficiency, suggesting its suitability for applications where time is a critical constraint. For the best performing model, it is further illustrated in Fig. 10
(a) 200 Iterations
(b) 400 Iterations
(c) 800 Iterations
(d) 5000 Iterations
(e) 10000 Iterations
(f) 30000 Iterations
Figure 10. Precision Over Training Iterations for NeRFacto - Scenario 1
Overall Performance and Suitability:
In sum, NeRFacto emerges as the most robust model in terms of precision, recall, F1 score, and image quality metrics (PSNR, SSIM, LPIPS), making it highly suitable for applications demanding high accuracy and completeness in 3D modeling. However, its time inefficiency at higher iterations might restrict its use in time-sensitive contexts. Instant-NGP presents a good balance between performance and efficiency, making it a viable option for moderately demanding scenarios. Detailed results are given in Table 3, after complete training. For more granular look, consult the supplementary Table 6. The Precision-Recall Curves based on varying distance threshold after maximum training of 30,000 iterations is given in Fig. 11
Table 3. Performance metrics of NeRFs reconstruction techniques - Scenario I
Model Name Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time(s) $\downarrow$
Instant-NGP 24.66 90.62 38.77 23.41 0.81 0.17 756
NeRFacto 73.57 94.72 82.81 22.24 0.73 0.12 1938
TensoRF 9.58 43.34 15.69 14.69 0.55 0.66 1973
# of points (%)
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Meters
  Precision    Recall
Figure 11. Precision-Recall Curves (Scenario-I) based on varying distance threshold after 30,000 iterations
Insight 1: Computational Cost and Accuracy Trade-off in Instant-NGP and NeRFacto:
The steep increase in performance metrics with the number of iterations for both Instant-NGP and NeRFacto suggests that these models require a substantial amount of data processing to achieve high accuracy, which is critical in high-fidelity 3D modeling. However, this also implies a higher computational cost, which needs to be considered in practical applications.
Insight 2: Model Suitability in High-Detail 3D Reconstructions:
The significant disparity in the performance of TensoRF compared to the other two models, particularly in precision and recall, indicates that not all NeRF models are equally suited for tasks requiring high-detail 3D reconstructions. This highlights the importance of model selection based on the specific requirements of the application.
Insight 3: Divergence in 2D Image Quality and 3D Reconstruction in Instant-NGP:
A detailed examination reveals that Instant-NGP demonstrates notable strength in 2D image quality metrics such as PSNR, SSIM, and LPIPS, reflecting its ability to produce superior rendered image quality. However, this excellence in 2D imaging does not correspondingly extend to 3D reconstruction metrics like Precision, Recall, and F1 Score. This observation highlights a significant distinction in the challenges associated with optimizing for high-quality image rendering as opposed to achieving accurate 3D representations. The model’s adeptness at rendering highly detailed 2D images does not necessarily imply its effectiveness in accurately reconstructing complex 3D structures, particularly in the context of intricate plant models. This insight underscores the need for a nuanced approach in evaluating the performance of models that are tasked with both 2D image rendering and 3D spatial reconstruction.
3.2. Scenario II - Multiple Plants Indoors
The estimated camera poses from colmap are visualized in Fig. 12. We observe marked differences in model behaviors compared to the single plant scenario, which are likely attributed to the added intricacy of multiple plants in a single scene. Results are illustrated in the Fig. 13.
Figure 12. Camera Pose Estimations for Scenario-II
(a) PSNR
(b) SSIM
(c) LPIPS
(d) Precision
(e) Recall
(f) F1 Score
Number of Iterations
  Instant-NGP    NeRFacto    TensoRF
Figure 13. Combined Comparison of 2D Image Quality (top) and 3D Geometry (bottom) Metrics for Scenario-II
Precision: Instant-NGP shows a steady increase in precision with more iterations, starting from 0.57 at 100 iterations and peaking at 23.45 at 30000 iterations. However, NeRFacto starts at a higher precision of 7.77 at 100 iterations and reaches a higher peak of 64.47 at 30000 iterations, indicating a more accurate reconstruction of the corn plants. TensoRF, although it improves with more iterations, lags behind the others, starting at 1.51 and reaching 20.50 at 30000 iterations. It is further illustrated in Fig. 14.
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Figure 14. Precision of 3D reconstruction of NeRF techniques - Scenario II
Recall:
For recall, a similar pattern is observed. Instant-NGP’s recall increases from 1.23 to 58.57, NeRFacto from 27.44 to 76.80, and TensoRF from 12.62 to 55.34 across iterations. NeRFacto consistently maintains a higher recall, suggesting its superior ability to encompass points in the ground truth.. It is further illustrated in Fig. 15.
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Figure 15. Recall of 3D reconstruction of NeRF techniques - Scenario II
F1 Score:
The F1 Score, balancing precision and recall, also follows this trend. Instant-NGP’s F1 score peaks at 33.49, NeRFacto at 70.10, and TensoRF at 29.91, all at 30000 iterations. NeRFacto demonstrates the best balance between precision and recall.
PSNR:
In terms of PSNR, which evaluates the quality of rendered images, all models show improvement with more iterations. Instant-NGP goes from 13.70 to 19.08, NeRFacto from 14.93 to 18.93, and TensoRF from 13.38 to 15.54. Instant-NGP achieves the highest PSNR, suggesting better image quality.
SSIM:
For SSIM, higher values indicate better image structure similarity. Instant-NGP progresses from 0.36 to 0.64, NeRFacto from 0.35 to 0.64, and TensoRF from 0.35 to 0.42. Both Instant-NGP and NeRFacto perform similarly and better than TensoRF in this aspect.
LPIPS:
Lower LPIPS values signify higher perceptual similarity to the ground truth. Instant-NGP decreases from 0.89 to 0.31, NeRFacto from 0.77 to 0.25, and TensoRF from 0.83 to 0.56, with NeRFacto showing the best perceptual image quality.
Computation Time:
The time taken for iterations is crucial for efficiency. Instant-NGP and NeRFacto have comparable times, but TensoRF takes significantly longer at higher iterations, indicating less time efficiency.
Overall Performance and Suitability:
NeRFacto emerges as the most balanced and efficient model, exhibiting high precision, recall, and F1 scores, along with favorable PSNR, SSIM, and LPIPS values. Its efficiency in time taken is also comparable to Instant-NGP. Instant-NGP, while showing improvements, doesn’t quite match NeRFacto’s balance of precision and recall. TensoRF, despite its merits, falls behind in several key metrics, particularly in precision, recall, SSIM, and LPIPS. Detailed results are given in Table 4, after complete training. For more granular look, consult the supplementary Table 7. The Precision-Recall Curves based on varying distance threshold after maximum training of 30,000 iterations is given in Fig. 16.
Table 4. Performance metrics of NeRFs reconstruction techniques - Scenario II
Model Name Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time(s) $\downarrow$
Instant-NGP 23.45 58.57 33.49 19.08 0.64 0.31 1886
NeRFacto 64.47 76.8 70.1 18.93 0.64 0.25 1226
TensoRF 20.5 55.34 29.91 15.54 0.42 0.56 2607
# of points (%)
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Meters
  Precision    Recall
Figure 16. Precision-Recall Curves (Scenario-II) based on varying distance threshold after 30,000 iterations
Insight 1: Improved Performance of TensoRF in Scenario 2:
In the second scenario, TensoRF demonstrated a notable improvement compared to its performance in the first scenario. Specifically, its F1 score, a critical metric for 3D modeling accuracy, significantly increased from 15.69 in the first scenario to 29.91 after 30,000 iterations in the second scenario. This substantial improvement highlights TensoRF’s potential in more complex or demanding 3D modeling tasks, especially when allowed to complete its training process.
Insight 2: 2D Metrics Versus 3D F1 Score for Instant-NGP and NeRFacto:
While Instant-NGP and NeRFacto show comparable results in 2D image quality metrics such as PSNR and SSIM, a distinct difference is observed in their 3D modeling capabilities, as reflected in their F1 scores, as observed in last scenario. This suggests that NeRFacto might be a more reliable choice for applications requiring high accuracy in 3D reconstructions.
3.3. Scenario III - Multiple Plants Outdoors
The estimated camera poses from colmap are visualized in Fig. 17
Figure 17. Camera Pose Estimations for Scenario-III
The results from our evaluation of three state-of-the-art NeRF techniques for 3D plant structure reconstruction are illustrated in the Fig. 18. Please note that the value of threshold distance $d$ for 3d metrics was considered 0.01, as compared to 0.005 in last two scenarios.
(a) PSNR
(b) SSIM
(c) LPIPS
(d) Precision
(e) Recall
(f) F1 Score
Number of Iterations
  Instant-NGP    NeRFacto    TensoRF
Figure 18. Combined Comparison of 2D Image Quality (top) and 3D Geometry (bottom) Metrics for Scenario-III
Precision:
Across all iterations, NeRFacto consistently demonstrates the highest precision, peaking at 68.29% at 60000 iterations, suggesting its superior ability to reconstruct points close to the ground truth. Instant-NGP shows a steady increase in precision with more iterations, yet never surpasses NeRFacto, reaching a maximum of 15.06%. TensoRF, while starting lower, reaches a comparable precision to Instant-NGP at higher iterations (40.95% at 60000 iterations). It is further illustrated in Fig. 19.
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Figure 19. Precision of 3D reconstruction of NeRF techniques - Scenario II
Recall:
NeRFacto also leads in recall, achieving a high of 82.32% at 60000 iterations, indicating its effectiveness in encompassing points from the ground truth. Instant-NGP shows significant improvement with increased iterations, reaching a recall of 59.55%, but remains behind NeRFacto. TensoRF shows substantial growth in recall, ultimately achieving 75.62%, positioning it between Instant-NGP and NeRFacto in terms of completeness. It is further illustrated in Fig. 20.
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Figure 20. Recall of 3D reconstruction of NeRF techniques - Scenario II
F1 Score:
Reflecting the balance between precision and recall, the F1 scores show NeRFacto as the superior model, with a peak score of 74.65% at 60000 iterations. Instant-NGP’s F1 score also improves with more iterations but tops out at 24.04%, significantly lower than NeRFacto. TensoRF’s F1 score, while lower than NeRFacto’s, surpasses Instant-NGP, reaching 53.13% at 60000 iterations.
PSNR:
In terms of image quality, as measured by PSNR, all models show incremental improvements with more iterations. Instant-NGP and NeRFacto display similar trends, with NeRFacto slightly leading, peaking at 16.70 at 60000 iterations. TensoRF shows a comparable maximum PSNR of 17.32, suggesting its slight edge in rendering higher-quality images.
SSIM:
For the SSIM metric, all three models show improvements with increased iterations. NeRFacto maintains a slight advantage over the others, peaking at 0.32 at 60000 iterations, indicating its better performance in maintaining structural integrity in the rendered images. Instant-NGP and TensoRF show similar SSIM scores, with TensoRF slightly leading at higher iterations.
LPIPS:
The LPIPS scores, which assess perceptual similarity, decrease for all models with more iterations, indicating improved performance. NeRFacto and TensoRF show similar trends, with NeRFacto having a slight edge, achieving a score of 0.34 at 60000 iterations compared to TensoRF’s 0.55. Instant-NGP’s performance is consistently lower in this metric.
Computation Time:
In terms of efficiency, Instant-NGP and NeRFacto are the fastest, followed by TensoRF.
Overall Performance and Suitability:
NeRFacto again emerges as the most balanced and robust model, excelling in precision, recall, F1 score, and LPIPS. Detailed results are given in Table 5, after complete training. For more granular look, consult the supplementary Table 8. The Precision-Recall Curves based on varying distance threshold after maximum training of 60,000 iterations is given in Fig. 21.
Table 5. Performance metrics of NeRFs reconstruction techniques - Scenario III
Model Name Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time(s) $\downarrow$
Instant-NGP 15.06 59.55 24.04 18.54 0.47 0.4 1466
NeRFacto 68.29 82.32 74.65 16.7 0.32 0.34 1499
TensoRF 40.95 75.62 53.13 17.32 0.39 0.55 1965
# of points (%)
(a) Instant-NGP
(b) TensoRF
(c) NeRFacto
Meters
  Precision    Recall
Figure 21. Precision-Recall Curves (Scenario-III) based on varying distance threshold after 60,000 iterations
Insight 1: Enhanced Performance of TensoRF in Outdoor Settings:
TensoRF demonstrates significant improvement in its performance in the third scenario compared to the first. Specifically, its F1 score has seen a notable increase; from 15.69 in the first scenario to 29.91 in the second, and reaching 53.13 after 30,000 iterations in the current outdoor scenario. This upward trajectory in F1 scores, which is a balanced measure of precision and recall, indicates TensoRF’s enhanced capability in outdoor environments, potentially outperforming Instant-NGP in these settings. This suggests that TensoRF might be a more suitable choice for outdoor 3D modeling tasks where both precision and completeness are crucial. This property may have contributed in the selection of TensoRF as a building block for using multiple local radiance fields, during in-the-wild reconstruction, in the recent paper [Meuleman et al., 2023].
Insight 2: LPIPS as a Strong Indicator of 3D Model Quality:
The LPIPS metric appears to be a more representative measure of the quality of the resulting 3D models. In the analysis, we observe that models with lower LPIPS scores consistently show better performance across other metrics. This trend indicates the relevance of LPIPS in assessing the perceptual quality of 3D models. The further investigation into how LPIPS correlates with other metrics could provide deeper insights into model performance, especially in the context of realistic and perceptually accurate 3D reconstructions.
4. Discussion
The difference in the output quality between Instant-NGP and Nerfacto, especially concerning the density and crispness of the rendered scenes, could indeed be related to the sampling strategies used by each algorithm.
Instant-NGP Sampling Strategy:
Instant-NGP uses an improved training and rendering algorithm that involves a ray marching scheme with an occupancy grid. This means that when the algorithm shoots rays into the scene to sample colors and densities, it uses an occupancy grid to skip over empty space, as well as areas behind high-density regions to improve efficiency.
  • The occupancy grid used in Instant-NGP is a multiscale grid that coarsely marks empty and non-empty space and is used to determine where to skip samples to speed up processing.
  • This approach is quite effective in terms of speed, leading to significant improvements over naive sampling methods.
  • However, if the occupancy grid isn’t fine-grained enough or if the method for updating this grid isn’t capturing the scene’s density variations accurately, it could lead to a "muddy" or overly dense rendering because it might not be sampling the necessary areas with enough precision.
Nerfacto Sampling Strategy:
Nerfacto, on the other hand, uses a combination of different sampling techniques:
  • Camera Pose Refinement: By refining camera poses, Nerfacto ensures that the samples taken are based on more accurate viewpoints, which directly affects the clarity of the rendered images.
  • Piecewise Sampler: This sampler is used to produce an initial set of samples, with a distribution that allows both dense sampling near the camera and appropriate sampling further away. This could lead to clearer images since it captures details both near and far from the camera.
  • Proposal Sampler: This is a key part of the Nerfacto method. It uses a proposal network to concentrate sample locations in regions that contribute most to the final render, usually around the first surface intersection. This targeted sampling could be a major reason why Nerfacto produces crisper images—it focuses computational resources on the most visually significant parts of the scene.
  • Density Field: By using a density field guided by a hash encoding and a small fused MLP, Nerfacto can efficiently guide sampling even further. It doesn’t require an extremely detailed density map since it is used primarily for guiding the sampling process, which means that it balances quality and speed without necessarily impacting the final image’s detail.
Comparing Sampling Strategies:
Instant-NGP’s sampling strategy is built for speed, with an occupancy grid that helps skip irrelevant samples. This approach is great for real-time applications but can potentially miss subtle density variations, leading to a denser and less clear output if the grid isn’t capturing all the necessary detail. Nerfacto’s sampling strategy is more complex and layered, with multiple mechanisms in place to ensure that sampling is done more effectively in areas that greatly affect the visual output. The combination of pose refinement, piecewise sampling, proposal sampling, and an efficient density field leads to more accurate sampling, which in turn produces crisper images. In summary, the reason for Nerfacto’s superior crispness likely stems from its more refined and targeted approach to sampling, which concentrates computational efforts on the most visually impactful parts of the scene. In contrast, Instant-NGP’s faster but less targeted sampling may result in less clarity and more visual artifacts.
5. Conclusions
The findings of this research are important for precision agriculture, offering a non-destructive and efficient method for detailed plant analysis. The ability to accurately reconstruct plant structures in 3D without physically altering or harming the plant is a key advancement. This non-intrusive approach enables essential agricultural tasks such as growth monitoring, yield prediction, and early disease detection to be performed more effectively, without the need for physical sampling or destruction of the plants. Furthermore, this research extends its impact beyond immediate applications.
In addition to these methodological advancements, our study contributes valuable resources for ongoing research. We collected a comprehensive dataset comprising images and ground truth 3D LiDAR scans, which will be instrumental for further studies. This dataset not only facilitates the testing and validation of current models but also serves as a benchmark for the development and evaluation of future NeRF models in agricultural contexts. Moreover, we developed a robust framework for the evaluation of NeRF models, which can be adapted and utilized in future research. This framework is designed to assess various aspects of model performance, ensuring a comprehensive evaluation of new techniques as they emerge. Moreover, this groundwork sets the stage for our next focus.
Our research also provides detailed insights into the performance of each model, identifying their strengths and limitations in various scenarios. These insights and resources pave the way for future research at the nexus of NeRF technologies and agriculture. The detailed analysis and case studies presented in this paper will be invaluable for researchers and practitioners looking to leverage NeRF in agricultural applications, contributing to the development of more advanced, efficient, and sustainable agricultural practices.
Supplement
Table 6. Performance metrics of NeRFs reconstruction techniques - Scenario I
Model Name Iterations Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time(s) $\downarrow$
Instant-NGP 100 0.29 2.48 0.53 17.14 0.58 0.81 13
Instant-NGP 200 0.32 1.50 0.53 18.23 0.56 0.75 26
Instant-NGP 400 1.25 7.14 2.13 19.64 0.60 0.66 42
Instant-NGP 800 5.66 35.17 9.74 21.21 0.64 0.55 61
Instant-NGP 1000 3.93 27.61 6.88 20.71 0.57 0.59 76
Instant-NGP 5000 21.93 89.03 35.20 22.73 0.76 0.28 175
Instant-NGP 10000 25.98 92.59 40.57 23.20 0.79 0.22 297
Instant-NGP 20000 23.21 88.38 36.77 23.42 0.81 0.18 527
Instant-NGP 30000 24.66 90.62 38.77 23.41 0.81 0.17 756
NeRFacto 100 1.94 20.98 3.55 18.11 0.59 0.75 14
NeRFacto 200 7.72 42.84 13.08 19.50 0.57 0.65 27
NeRFacto 400 20.86 68.64 32.00 21.27 0.64 0.55 43
NeRFacto 800 39.48 80.26 52.92 22.33 0.67 0.45 64
NeRFacto 1000 41.35 82.54 55.09 22.20 0.66 0.44 79
NeRFacto 5000 66.43 92.51 77.33 22.30 0.73 0.19 430
NeRFacto 10000 70.04 93.94 80.25 22.34 0.74 0.15 564
NeRFacto 20000 73.32 94.51 82.58 22.35 0.74 0.13 1068
NeRFacto 30000 73.57 94.72 82.81 22.24 0.73 0.12 1938
TensoRF 100 0.43 2.27 0.72 13.44 0.54 0.82 13
TensoRF 200 0.65 4.47 1.13 13.55 0.52 0.82 25
TensoRF 400 1.18 8.76 2.07 13.51 0.51 0.81 40
TensoRF 800 1.86 14.22 3.29 13.10 0.50 0.79 61
TensoRF 1000 2.05 16.40 3.65 13.14 0.49 0.79 77
TensoRF 5000 6.63 36.57 11.22 13.59 0.52 0.70 420
TensoRF 10000 9.58 43.47 15.69 14.64 0.55 0.67 859
TensoRF 20000 9.51 43.19 15.59 14.68 0.55 0.67 1651
TensoRF 30000 9.58 43.34 15.69 14.69 0.55 0.66 1973
Table 7. Performance metrics of NeRFs reconstruction techniques - Scenario II
Model Name Iterations Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time(s) $\downarrow$
Instant-NGP 100 0.57 1.23 0.78 13.70 0.36 0.89 13
Instant-NGP 200 1.52 4.41 2.27 13.91 0.32 0.85 27
Instant-NGP 400 2.49 9.35 3.93 14.95 0.36 0.77 43
Instant-NGP 800 5.83 24.29 9.41 15.91 0.39 0.69 63
Instant-NGP 1000 5.72 22.80 9.14 15.68 0.32 0.69 78
Instant-NGP 5000 18.30 51.68 27.02 18.12 0.55 0.44 176
Instant-NGP 10000 20.86 55.25 30.28 18.67 0.59 0.38 296
Instant-NGP 20000 23.24 58.53 33.27 19.05 0.63 0.33 1023
Instant-NGP 30000 23.45 58.57 33.49 19.08 0.64 0.31 1886
NeRFacto 100 7.77 27.44 12.12 14.93 0.35 0.77 15
NeRFacto 200 16.42 37.77 22.89 15.51 0.32 0.70 28
NeRFacto 400 25.83 49.81 34.02 16.40 0.39 0.67 44
NeRFacto 800 35.97 56.86 44.07 17.10 0.43 0.60 64
NeRFacto 1000 36.22 57.80 44.53 16.96 0.40 0.61 79
NeRFacto 5000 58.64 70.87 64.18 18.69 0.59 0.37 404
NeRFacto 10000 61.31 74.12 67.11 18.91 0.62 0.31 749
NeRFacto 20000 63.68 76.21 69.38 19.00 0.64 0.27 988
NeRFacto 30000 64.47 76.80 70.10 18.93 0.64 0.25 1226
TensoRF 100 1.51 12.62 2.70 13.38 0.35 0.83 14
TensoRF 200 2.53 17.53 4.43 13.81 0.34 0.79 27
TensoRF 400 4.41 26.94 7.58 14.19 0.34 0.74 44
TensoRF 800 6.28 35.10 10.66 14.46 0.34 0.71 66
TensoRF 1000 7.09 35.61 11.82 14.51 0.33 0.70 82
TensoRF 5000 13.09 51.34 20.86 15.16 0.36 0.64 399
TensoRF 10000 20.51 55.03 29.88 15.51 0.42 0.57 840
TensoRF 20000 20.44 54.97 29.80 15.54 0.42 0.56 1709
TensoRF 30000 20.50 55.34 29.91 15.54 0.42 0.56 2607
Table 8. Performance metrics of NeRFs reconstruction techniques - Scenario III
Model Name Iterations Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time(s) $\downarrow$
Instant-NGP 100 1.61 17.61 2.95 15.31 0.29 0.91 16
Instant-NGP 200 1.29 8.19 2.23 14.66 0.27 0.84 30
Instant-NGP 400 2.45 19.22 4.34 15.65 0.29 0.82 46
Instant-NGP 800 2.81 18.90 4.89 15.91 0.29 0.79 67
Instant-NGP 1000 2.88 10.08 4.48 15.61 0.26 0.71 82
Instant-NGP 5000 6.47 40.28 11.15 17.39 0.37 0.61 186
Instant-NGP 10000 9.38 47.12 15.65 17.86 0.40 0.54 311
Instant-NGP 20000 11.62 54.26 19.13 18.27 0.43 0.48 548
Instant-NGP 30000 12.96 56.88 21.11 18.46 0.45 0.45 783
Instant-NGP 60000 15.06 59.55 24.04 18.54 0.47 0.40 1466
NeRFacto 100 4.26 35.99 7.61 15.65 0.29 0.84 13
NeRFacto 200 6.53 42.49 11.31 15.91 0.27 0.66 27
NeRFacto 400 10.48 50.45 17.35 16.44 0.29 0.69 43
NeRFacto 800 18.49 57.01 27.93 16.81 0.31 0.69 64
NeRFacto 1000 18.39 58.02 27.93 16.56 0.29 0.65 80
NeRFacto 5000 49.87 74.75 59.83 17.29 0.34 0.54 189
NeRFacto 10000 58.22 78.80 66.96 17.24 0.34 0.47 318
NeRFacto 20000 64.91 80.61 71.92 17.03 0.34 0.41 561
NeRFacto 30000 66.30 81.33 73.05 16.88 0.33 0.38 803
NeRFacto 60000 68.29 82.32 74.65 16.70 0.32 0.34 1499
TensoRF 100 1.48 12.65 2.65 15.05 0.29 0.96 17
TensoRF 200 3.05 27.40 5.48 15.26 0.29 0.90 31
TensoRF 400 5.69 43.45 10.06 15.63 0.29 0.83 47
TensoRF 800 8.92 49.59 15.13 15.96 0.29 0.78 68
TensoRF 1000 10.48 50.82 17.37 16.03 0.29 0.75 84
TensoRF 5000 28.74 70.83 40.89 16.84 0.34 0.64 208
TensoRF 10000 40.85 75.24 52.95 17.31 0.38 0.56 374
TensoRF 20000 40.82 75.38 52.96 17.33 0.39 0.55 697
TensoRF 30000 40.80 75.26 52.92 17.34 0.39 0.55 1018
TensoRF 60000 40.95 75.62 53.13 17.32 0.39 0.55 1965
(a) 200 Iterations
(b) 400 Iterations
(c) 800 Iterations
(d) 5000 Iterations
(e) 10000 Iterations
(f) 30000 Iterations
Figure 22. Precision Over Training Iterations for NeRFacto - Scenario II
Table 9. Performance metrics of MipNeRF (failed) reconstruction - Scenario II
Iterations Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (s) $\downarrow$
100 0.47 7.57 0.89 11.01 0.34 0.91 26
200 0.63 10.67 1.19 11.24 0.34 0.93 52
400 0.69 12.1 1.3 11.07 0.34 0.92 86
800 0.48 7.68 0.91 9.88 0.33 0.94 158
1000 0.44 7.19 0.82 9.65 0.32 0.94 192
5000 0.35 5.41 0.66 9.31 0.31 0.94 687
10000 0.42 6.29 0.79 8.88 0.29 0.9 1303
20000 0.34 5.3 0.65 8.78 0.28 0.89 2508
30000 0.35 5.71 0.67 9 0.27 0.88 3856